OpenStudy (anonymous):

Solve the equation. 16y^2+8y+1=0

OpenStudy (anonymous):

Use the quadratic equation:\[\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (lovelycharm):

Simplifying 16y2 + 8y + 1 = 0 Reorder the terms: 1 + 8y + 16y2 = 0 Solving 1 + 8y + 16y2 = 0 Solving for variable 'y'. Factor a trinomial. (1 + 4y)(1 + 4y) = 0 Subproblem 1Set the factor '(1 + 4y)' equal to zero and attempt to solve: Simplifying 1 + 4y = 0 Solving 1 + 4y = 0 Move all terms containing y to the left, all other terms to the right. Add '-1' to each side of the equation. 1 + -1 + 4y = 0 + -1 Combine like terms: 1 + -1 = 0 0 + 4y = 0 + -1 4y = 0 + -1 Combine like terms: 0 + -1 = -1 4y = -1 Divide each side by '4'. y = -0.25 Simplifying y = -0.25 Subproblem 2Set the factor '(1 + 4y)' equal to zero and attempt to solve: Simplifying 1 + 4y = 0 Solving 1 + 4y = 0 Move all terms containing y to the left, all other terms to the right. Add '-1' to each side of the equation. 1 + -1 + 4y = 0 + -1 Combine like terms: 1 + -1 = 0 0 + 4y = 0 + -1 4y = 0 + -1 Combine like terms: 0 + -1 = -1 4y = -1 Divide each side by '4'. y = -0.25 Simplifying y = -0.25Solutiony = {-0.25, -0.25}